# Safe and Efficient In-Context Learning via Risk Control

### v1

@@ Abstract
Large language models (LLMs) demonstrate a remarkable ability to learn new tasks from a few in-context examples. However, this flexibility introduces safety concerns: LLMs can be influenced by incorrect or malicious demonstrations – for example, if an adversary tampers with or injects harmful examples without a human supervisor noticing. This motivates principled designs in which the system itself includes built-in mechanisms to guard against such attacks. We propose a novel approach to limit the degree to which harmful demonstrations can degrade model performance. First, we define a baseline “safe” behavior for the model – the model’s performance given no in-context demonstrations (zero-shot). Next, we apply distribution-free risk control (DFRC) to control the extent to which in-context samples can decay performance below zero-shot. We achieve this by leveraging dynamic early exit prediction, ignoring later attention heads that attend the most to the unsafe inputs. Finally, we propose modifications to DFRC that allow it to both control risk for harmful inputs and leverage performance and efficiency gains on helpful inputs. We present both theoretical and empirical results showing that our approach can effectively control risk for harmful in-context demonstrations while simultaneously achieving substantial computational efficiency gains with helpful demonstrations.
@@ 1 Introduction
Large language models (LLMs) have shown an impressive ability to be adapted to a wide variety of tasks through methods such as prompt tuning and in-context learning, many of which require only minimal data and do not require expensive fine-tuning. Yet this adaptability introduces safety concerns: incorrect, adversarial, or otherwise harmful demonstrations can degrade performance or elicit unsafe outputs. Imagine an LLM deployed for a specific use case and adapted through in-context demonstrations; these demonstrations may be misleading for a number of reasons – such as unintended user error or intentional tampering by an adversary (such as many-shot jailbreaking (Anil et al., 2024) and prompt injections (Liu et al., 2024; Das et al., 2024)). For instance, consider a LLM coding agent prompted with example API calls. A developer may mistakenly provide hardcoded credentials, accidentally teaching the model to replicate insecure code; an adversary could insert a demonstration where user validation is bypassed, introducing security vulnerabilities. Such vulnerabilities could escape the notice of a human system designer, motivating us to develop built-in safeguards so that the LLM defaults to disregarding these compromised demonstrations.
@@
In this paper, we apply distribution-free risk control (DFRC) to mitigate the influence of corrupted in-context examples by comparing the loss of the adapted model to the default zero-shot model. LLMs’ zero-shot performance on a wide variety of tasks has become quite strong in recent years and continues to improve (Kojima et al., 2023). This makes zero-shot LLMs comparatively well-understood and predictable, whereas ICL models on arbitrary user-supplied demonstrations may reflect uncontrolled or adversarial distribution shifts. Using the zero-shot model as a baseline also anchors risk control in a setting that has undergone extensive pre-deployment safety testing (Zhang et al., 2024; Yuan et al., 2025), unlike the highly variable in-context examples that arise during deployment. Recent work has shown that depth controls how much an LLM can learn from in-context examples; LLMs “overthink” on harmful examples, meaning their performance peaks at some intermediate layer and drops in deeper (later) layers (Halawi et al., 2024). We replicate these results on our tasks; an example is shown in Fig. 1(b), where the model does well given correct demonstrations, reasonably well given no demonstrations (zero-shot), and very poorly given incorrect demonstrations. Inspired by this, we implement early-exiting as a natural mechanism for applying this risk control.
@@
In summary, our contributions are as follows: We propose three important contributions to enable safety for ICL: a novel formulation of early-exit models for safety using the safe zero-shot baseline (§3.1), a novel ICL loss designed to measure overthinking (§3.3), and a simple adaptation of the Learn-then-Test (LTT) risk control framework that balances safety with efficiency gains (§3.4). We show, via extensive experiments, that our approach can effectively prevent overthinking while still allowing the LLM to benefit from helpful demonstrations, providing robust safety guarantees even with mixed-quality inputs (§4.1) and enabling major computational efficiency gains compared to prior approaches (§4.2). Experiments across 8 diverse benchmark tasks and 4 distinct models show that our framework is able to guarantee safety on model outputs relative to zero-shot, while simultaneously achieving a greater than 50% speedup in comparison to previous approaches. Overall, to the best of our knowledge, this is the first work to establish a principled framework that controls the risk of harmful in-context demonstrations, while simultaneously leveraging dynamic early exit mechanisms to achieve performance and computational efficiency gains with helpful demonstrations.
Figure 1: (a) A LLM is given in-context demonstrations of unknown quality (helpful or harmful). The model needs to infer whether to rely on the answer it obtains using the given demonstrations without knowing ahead of time if they are helpful or not. If not, it falls back to the answer it would give without seeing any demonstrations at all (zero-shot). (b) When given incorrect demonstrations, it is better to either early-exit or simply not use the given demonstrations than to use the model’s final prediction – staying in the “safe” performance range between zero-shot and correct demonstrations.
@@ 2 Preliminaries
Data
Let $x\in\mathcal{X}$ denote an input text (e.g., a question), and let $y\in\mathcal{Y}$ be its associated label (e.g., a choice from a predetermined set of possible answers). Given our focus on classification, the label space is defined as $\mathcal{Y}=\{1,\ldots,K\}$, with $K$ representing the number of possible answers. Moreover, we denote a context set of $N_{c}$ demonstrations as $c=\{(x_{i},y_{i})\}_{i=1}^{N_{c}}$. Lastly, a data-generating distribution over $\mathcal{X}\times\mathcal{Y}$ is denoted with $\mathcal{P}$.
Model
We denote by $p(\cdot|x,c)$ an LLM that takes a given input $x$ together with the context set $c$, and outputs a probability distribution over possible labels $k\in\mathcal{Y}$. We sample both the input $x$ and the context $c$ from the same dataset, but explicitly prevent any overlap between $x$ and the elements of $c$ for a given prompt. By excluding $x$ from $c$, we can prevent the model “copying” answers from the context rather than demonstrating meaningful in-context learning behavior, while still drawing both from the same distribution to avoid any bias in context/input dataset selection.
2.1 Early-Exit Language Models for Classification Tasks
Traditionally, LLMs pass through all $L$ layers of the model before making a prediction. In contrast, early-exit LLMs (Elbayad et al., 2020; Schuster et al., 2022) offer the option to yield a prediction after each layer. This is achieved by passing the current hidden representation through an “unembedding” matrix that maps the hidden state to the vocabulary, pruned to the finite set of possible labels for a particular task. Specifically, for each layer $l\in\{1,\ldots,L\}$ where $L$ represents the total number of layers in the model, a confidence score $\mathfrak{C}_{l}\in[0,1]$ and an exit threshold $\lambda\in[0,1]$ are defined. An early prediction is returned as soon as the confidence at the current layer exceeds the threshold, with each condition evaluated in the order of the layers in the original model. This is outlined in Eq. 1 and illustrated in Fig. 2.
@@ 2.2 Controlling Predictive Risk via Distribution-Free Risk Control
Figure 3: Some choices of $\lambda$ thresholds can both attain performance gains from correct demonstrations and control overthinking from incorrect demonstrations. The highlighted regions show where $\lambda$ values exist such that we lose no more than 5% of the accuracy gains from correct demonstrations while still doing better than the full model given incorrect demonstrations.
3 In-Context Learning Risk Control via Early-Exit
In this section, we detail our approach to mitigating overthinking for in-context learning, i.e., preventing the LLM from picking up on harmful demonstrations by combining early-exiting with risk control. We note that across all our tasks and datasets, correct in-context demonstrations improve performance above zero-shot, and incorrect demonstrations harm performance to worse than zero-shot; so we use the terms “correct” vs “helpful” and “incorrect” vs “harmful” interchangeably.
We begin by introducing a novel formulation of early-exit models for safety by using the safe zero-shot baseline (§3.1). We then show that applying our early-exit approach can effectively prevent overthinking while still allowing the LLM to benefit from helpful demonstrations, provided an appropriate exit threshold $\lambda$ is chosen (§3.2). Next, we propose a novel ICL loss designed to measure overthinking on which we can apply risk control to ensure safety on new test points (§3.3). Finally, we introduce a simple adaptation of the Learn-then-Test (LTT) framework to accommodate losses that may take on negative values, as is often the case in ICL (§3.4).
3.1 Safe In-Context Learning Predictor
We begin with our base in-context learning model $p$ given input $x$ and in-context demonstrations $c$ without early-exit, $p(\cdot|x,c)$. This model can “overthink” (overfit to potentially harmful demonstrations $c$), significantly degrading the output over the zero-shot model, $p(\cdot|x)$. We propose to define a new safe in-context learning model $\bar{p}_{\lambda}(\cdot|x,c)$ by augmenting this base model for in-context learning in the following two ways: (i) enable the model to make predictions from intermediate exits given a confidence threshold and (ii) if at no exit the confidence exceeds the threshold, ignore the context $c$ and use the zero-shot prediction. With these augmentations, the safe ICL model can be defined as:
$\bar{y}_{\lambda}:=\begin{cases}\arg\max\limits_{k\in\mathcal{Y}}p_{1}\left(k|x,c\right)&\text{if }\mathfrak{C}_{1}\geq\lambda,\\ \vdots&\vdots\\ \arg\max\limits_{k\in\mathcal{Y}}p_{L}\left(k|x,c\right)&\text{else if }\mathfrak{C}_{L}\geq\lambda,\\ \mathbf{\arg\max\limits_{k\in\mathcal{Y}}p_{L}\left(k|x\right)}&\textbf{otherwise.}\end{cases}$ (2)
Note that such a predictor enables us to leverage early-exit to gain efficiency and performance on helpful demonstrations, while also intervening early to avoid harmful demonstrations before the model fully processes them. When early-exit alone is insufficient to guarantee safety, the zero-shot model serves as a reliable baseline (see the last condition in Eq.2).
@@ 3.2 Early Exit Reduces Overthinking
As observed by Halawi et al. (2024), overthinking primarily arises in the deeper (i.e., later) layers of models, where it can override the strong inductive biases or correct predictions established in earlier layers. This phenomenon can result in degraded or unsafe outputs, as illustrated in Fig. 1. While prior work proposes pruning certain attention heads in the later layers as a preventive measure (Halawi et al., 2024), we instead advocate for leveraging dynamic early exiting (Teerapittayanon et al., 2016) which has previously been employed to mitigate overthinking in settings outside the scope of LLMs and in-context learning (Kaya et al., 2019; Jazbec et al., 2023). Notably, early exiting offers a natural solution to the overthinking problem: by terminating inference at an intermediate layer, it prevents the model from fully processing potentially misleading context. Thus, stopping early increases accuracy when the context is harmful. We show that overthinking behavior occurs on all our tasks in Fig. 17. Although not our primary goal, one nice side benefit of early-exiting is efficiency gains, as not all layers are processed before outputting a prediction.
@@
We demonstrate the effectiveness of early exiting in mitigating ICL overthinking in Fig.3. Across most of the datasets considered, and for a broad range of exit thresholds $\lambda$, an early-exit LLM ($p_{\lambda}$) outperforms the full model ($p_{L}$) on inputs with incorrect demonstrations (§I). Importantly, for certain thresholds – the yellow highlighted area in Fig.3 – early termination also does not significantly degrade performance on samples with helpful demonstrations. These results underscore the potential of dynamic inference to prevent LLMs from picking up on harmful demonstrations.
3.3 In-Context Learning Risk
@@
After describing how early exiting can be used to prevent overthinking, we now turn to the problem of selecting an appropriate early-exit strategy using risk control, as introduced in §2.2. As a first step, we propose a novel in-context learning loss:
$\ell_{\text{ICL}}(\lambda;x,y,c):=\ell(\bar{y}_{\lambda}(x,c),y)-\ell(\hat{y}(x),y)\>,$ (3)
@@
where $\bar{y}_{\lambda}(x,c):=\arg\max_{k}\bar{p}_{\lambda}(k,x,c)$ and $\hat{y}(x):=\arg\max_{k}p_{L}(k|x)$ denote the predictions of the safe ICL model with demonstrations $c$ (see Eq. 2) and the full zero-shot model, respectively. Crucially, both predictions are produced by the same underlying LLM. The $\ell_{\text{ICL}}$ loss compares the performance of the early-exit model with demonstrations to that of the model without demonstrations. This formulation makes it well-suited for measuring overthinking: if the demonstrations are harmful, the loss will be positive; if they are helpful, the loss will be negative. In contrast, prior early-exiting work (Schuster et al., 2022; Jazbec et al., 2024) has focused exclusively on loss definitions that compare early-exit outputs to final outputs given the same input. Such losses are less effective for identifying and addressing overthinking, as they do not account for the drop in performance in later layers due to harmful in-context demonstrations. Our model thus allows for robust measurement of risk considering both correct and incorrect demonstrations, enabling effective risk control with mixed-quality context.
@@
$R_{\text{ICL}}(\hat{\lambda})=\mathbb{E}_{(x,y,c)\sim\mathcal{P}}[\ell_{\text{ICL}}(\hat{\lambda};x,y,c)]\leq\epsilon\>,$
where $\epsilon>0$ is a user-specified tolerance level representing the acceptable degree of overthinking. Naturally, smaller values of $\epsilon$ impose stricter control, prioritizing thresholds that suppress overthinking more aggressively. However, this may come at the cost of reduced performance on helpful demonstrations—a tradeoff we explore in §H. Since we observe that risks computed using the ICL loss $\ell_{\text{ICL}}$ are not monotonic with respect to $\lambda$ (Fig. 7), the Learn-then-Test (LTT) framework (Angelopoulos et al., 2021) is the only viable option; hence, we use LTT as our risk-control approach for selecting our exit threshold $\hat{\lambda}$.

### v2

# Controlling the Risk of Corrupted Contexts for Language Models via Early-Exiting

@@ Abstract
Large language models (LLMs) can be influenced by harmful or irrelevant context, which can significantly harm model performance on downstream tasks. This motivates principled designs in which LLM systems include built-in mechanisms to guard against such “garbage in, garbage out” scenarios. We propose a novel approach to limit the degree to which harmful context can degrade model performance. First, we define a baseline “safe” behavior for the model – the model’s performance given no context at all (zero-shot). Next, we apply distribution-free risk control (DFRC) to control the extent to which the user-provided context can decay performance below this safe zero-shot baseline. We achieve this by leveraging dynamic early exit prediction, ignoring later attention heads that attend the most to the unsafe inputs. Finally, we propose modifications to DFRC that allow it to both control risk for harmful inputs and leverage performance and efficiency gains on helpful inputs. We present both theoretical and empirical results across 9 tasks spanning in-context learning and open-ended question answering, showing that our approach can effectively control risk for harmful context and simultaneously achieve substantial computational efficiency gains with helpful context.
Machine Learning, Risk Control, Language Modeling, Early Exit Language Models, AI Safety
Figure 1: (a) A LLM is given some input context C of unknown quality (helpful or harmful). For a test question x, the model needs to infer whether to rely on the answer it obtains using the given context without knowing ahead of time if it was helpful or not. If not, it falls back to the answer it would give without seeing any additional context (zero-shot). (b) When given harmful context (such as incorrect in-context demonstrations), it is better to either early-exit or simply ignore the given context than to use the model’s final prediction – staying in the “safe” performance range between zero-shot and correct (or helpful) context.
@@ 1 Introduction
Large language models (LLMs) have shown an impressive ability to be adapted to a wide variety of tasks through methods such as prompt tuning and in-context learning, many of which require only minimal data and do not require expensive fine-tuning as they only require manipulating the context provided to the model. Yet this adaptability introduces safety concerns: incorrect, misleading, or otherwise harmful context can degrade performance or elicit unsafe outputs. Imagine an LLM deployed for a specific use case and adapted through user-provided context; this context may be misleading for a number of reasons – such as unintended user error, malicious intent, or task mis-specification. For instance, consider an LLM assisting clinicians with summarizing patient notes. A busy nurse might accidentally provide an incorrectly labeled demonstration, marking a non-urgent case as urgent due to time pressure. Alternatively, a biased clinician may omit or skew important details due to their perception about a particular patient. These kinds of inconsistent, low-quality context can shift the model’s behavior toward unreliable triage recommendations. Such errors could escape the notice of a human system designer, motivating us to develop built-in safeguards so that the LLM defaults to disregarding these compromised inputs.
@@
In this paper, we apply distribution-free risk control (DFRC) to mitigate the influence of corrupted context by comparing the loss of the adapted model to the zero-shot model. LLMs’ zero-shot performance on a wide variety of tasks has become quite strong in recent years and continues to improve (Kojima et al., 2023). This makes zero-shot LLMs comparatively well-understood and predictable, whereas in-context learning (ICL) or context-influenced models on arbitrary user-supplied inputs may reflect uncontrolled or adversarial distribution shifts. Using the zero-shot model as a baseline also anchors risk control in a setting that has undergone extensive pre-deployment safety testing (Zhang et al., 2024; Yuan et al., 2025), unlike the highly variable user-provided contexts that arise during deployment. Recent work has shown that depth controls how much an LLM can learn from its context; LLMs “overthink” on harmful inputs, meaning their performance peaks at some intermediate layer and drops in deeper (later) layers (Halawi et al., 2024). We replicate these results on our tasks; an example is shown in Fig. 1(b), where the model does well given helpful context, reasonably well given no context (zero-shot), and very poorly given harmful context. Inspired by this, we implement early-exiting as a natural mechanism for applying risk control.
@@
In summary, we propose three important contributions to ensure safety under arbitrary context: a novel formulation of early-exit models for safety (§3.1), a novel context-aware loss designed to measure overthinking (§3.3), and an important adaptation of the Learn-then-Test (LTT) risk control framework balancing safety with efficiency (§3.4). We show via extensive experiments that our approach prevents overthinking while still allowing the LLM to benefit from helpful context, providing robust safety guarantees even with mixed-quality inputs (§4.3) and enabling major computational efficiency gains compared to prior approaches (§4.4). We conduct experiments across 9 diverse tasks and 5 distinct models, spanning a wide variety of in-context learning settings and an open-ended, multi-token question answering task. Our framework guarantees safety on model outputs relative to zero-shot while simultaneously achieving a greater than 50% speedup in comparison to previous approaches across all tasks. Overall, to the best of our knowledge, this is the first work to establish a principled framework controlling the risk of potentially harmful user-provided context, while simultaneously leveraging dynamic early exit mechanisms to achieve performance and computational efficiency gains with helpful context.
@@ 2 Preliminaries
Data Let $x\in\mathcal{X}$ denote an input text (e.g., a question), and let $y\in\mathcal{Y}$ be its associated label (e.g., the associated answer). Given our focus on answer prediction, the label space is denoted by the set $\mathcal{Y}$, which may represent a finite set of valid classes (for classification) or an unbounded set of answers to a question. In the classification setting, $\mathcal{Y}$ is finite and can be defined as $\mathcal{Y}=\{1,\ldots,K\}$, with $K$ representing the number of possible answers. Moreover, we denote the context provided by the user to the model as $c$. In the case of in-context learning, $c$ can also be represented as a context set of $N_{c}$ demonstrations as $c_{\text{ICL}}=\{(x_{i},y_{i})\}_{i=1}^{N_{c}}$. Lastly, a data-generating distribution over $\mathcal{X}\times\mathcal{Y}$ is denoted with $\mathcal{P}$.
Model We denote by $p(\cdot|x,c)$ an LLM that takes a given input $x$ together with the context $c$ and outputs a probability distribution over possible labels in $\mathcal{Y}$. To construct a set of demonstrations $c$ for in-context learning, we sample both the input $x$ and the context set $c_{\text{ICL}}$ from the same dataset, but explicitly prevent any overlap between $x$ and the elements of $c_{\text{ICL}}$ for a given prompt. By excluding $x$ from $c_{\text{ICL}}$, we can prevent the model “copying” answers from the context rather than demonstrating meaningful in-context learning behavior, while still drawing both from the same distribution to avoid any bias in context/input dataset selection.
2.1 Early-Exiting for Classification Tasks
Traditionally, LLMs pass through all $L$ layers of the model before making a prediction. In contrast, early-exit LLMs (Elbayad et al., 2020; Schuster et al., 2022) offer the option to yield a prediction after each layer. This is achieved by passing the current hidden representation through an “unembedding” matrix that maps the hidden state to the vocabulary, pruned to the finite set of possible labels for classification tasks. Specifically, for each layer $l\in\{1,\ldots,L\}$ where $L$ represents the total number of layers in the model, a confidence score $\mathfrak{C}_{l}\in[0,1]$ and an exit threshold $\lambda\in[0,1]$ are defined. An early prediction is returned as soon as the confidence at the current layer exceeds the threshold, with each condition evaluated in the order of the layers in the original model, as outlined in Eq. 1 and illustrated in Fig. 2.
@@ 2.2 Controlling Predictive Risk via Distribution-Free Risk Control
Figure 3: Some choices of $\lambda$ thresholds can both attain performance gains from helpful context and control overthinking from harmful context. The highlighted regions show where $\lambda$ values exist such that we lose no more than 5% of the accuracy gains from correct (helpful) in-context demonstrations while still doing better than the full model given incorrect (harmful) in-context demonstrations.
3 Context-Aware Risk Control via Early-Exit
In this section, we detail our approach to mitigating overthinking under potentially harmful context, i.e., preventing the LLM from picking up on harmful context by combining early-exiting with risk control. We begin by introducing a novel formulation of early-exit models for safety by using the safe zero-shot model as a baseline (§3.1). We then show that applying our early-exit approach can effectively prevent overthinking while still allowing the LLM to benefit from helpful context, provided an appropriate exit threshold $\lambda$ is chosen (§3.2). Next, we propose a novel context-aware loss designed to measure overthinking, on which we can apply risk control to ensure safety on new test points (§3.3). Finally, we introduce a simple adaptation of the Learn-then-Test (LTT) framework to accommodate losses that may take on negative values, as is often the case when both helpful and harmful context is present in the data (§3.4).
3.1 Safe Context-Aware Predictor
We begin with our base model $p$ given input $x$ and context $c$ without early-exit, $p(\cdot|x,c)$. This model can “overthink” (overfit to potentially harmful context $c$), significantly degrading the output over the zero-shot model, $p(\cdot|x)$. We propose to define a new safe context-aware early exit model $\bar{p}_{\lambda}(\cdot|x,c)$ by augmenting this base model in the following two ways: (i) enable the model to make predictions from intermediate exits given a confidence threshold and (ii) if at no exit the confidence exceeds the threshold, ignore the context $c$ and use the zero-shot prediction. With these augmentations, the safe context-aware model can be defined as:
$\bar{y}_{\lambda}:=\begin{cases}\hat{y}_{\lambda}&\mathfrak{C}_{i}\geq\lambda,\ \exists i\in\{1,...,L\}\\ \mathbf{\arg\max\limits_{k\in\mathcal{Y}}p_{L}\left(k|x\right)}&\textbf{otherwise.}\end{cases}$ (2)
Note that such a predictor enables us to leverage early-exit to gain efficiency and performance on helpful context, while also intervening early to avoid harmful context before the model fully processes it. When early-exit alone is insufficient to guarantee safety, the zero-shot model serves as a reliable baseline (see the last condition in Eq.2).
@@ 3.2 Early Exit Reduces Overthinking
As observed by (Halawi et al., 2024), overthinking primarily arises in the deeper (i.e., later) layers of models, where it can override the strong inductive biases or correct predictions established in earlier layers. This phenomenon can result in degraded or unsafe outputs, as illustrated in Fig. 1. While prior work proposes pruning certain attention heads in the later layers as a preventive measure (Halawi et al., 2024), we instead advocate for leveraging dynamic early exiting (Teerapittayanon et al., 2016) which has previously been employed to mitigate overthinking in settings outside the scope of LLMs and in-context learning (Kaya et al., 2019; Jazbec et al., 2023). Notably, early exiting offers a natural solution to the overthinking problem: by terminating inference at an intermediate layer, it prevents the model from fully processing potentially misleading context. Thus, stopping early increases accuracy when the context is harmful. We show that overthinking behavior occurs on all our tasks in Fig. 21. Although not our primary goal, one nice side benefit of early-exiting is efficiency gains, as not all layers are processed before outputting a prediction.
@@
We demonstrate the effectiveness of early exiting in mitigating overthinking on harmful context in Fig.3. Across most of our datasets and for a broad range of exit thresholds $\lambda$, an early-exit LLM ($p_{\lambda}$) outperforms the full model ($p_{L}$) on inputs with harmful context (§I). Importantly, for certain thresholds – the yellow highlighted area in Fig.3 – early termination also does not significantly degrade performance on samples with helpful context. These results underscore the potential of dynamic inference to prevent LLMs from picking up on harmful context.
3.3 Context-Aware Prediction Risk
@@
After describing how early exiting can be used to prevent overthinking, we now turn to the problem of selecting an appropriate early-exit strategy using risk control, as introduced in §2.2. As a first step, we propose a novel context-aware prediction loss:
$\ell_{\text{c}}(\lambda;x,y,c):=\ell(\bar{y}_{\lambda}(x,c),y)-\ell(\hat{y}(x),y)\>,$ (3)
@@
where $\bar{y}_{\lambda}(x,c):=\arg\max_{k}\bar{p}_{\lambda}(k,x,c)$ and $\hat{y}(x):=\arg\max_{k}p_{L}(k|x)$ denote the predictions of the safe context-aware model with context $c$ (see Eq. 2) and the full zero-shot model, respectively. Crucially, both predictions are produced by the same underlying LLM. The $\ell_{\text{c}}$ loss compares the performance of the early-exit model with context to that of the model without context. This formulation makes it well-suited for measuring overthinking: if the context is harmful, the loss will be positive; if it is helpful, the loss will be negative. In contrast, prior early-exiting work (Schuster et al., 2022; Jazbec et al., 2024) has focused exclusively on loss definitions that compare early-exit outputs to final-layer outputs given the same input. Such losses are less effective for identifying and addressing overthinking, as they do not account for the drop in performance in later layers due to harmful context. Our model thus allows for robust measurement of risk considering both helpful and harmful context, enabling risk control with mixed context.
@@
$R_{\text{c}}(\hat{\lambda})=\mathbb{E}_{(x,y,c)\sim\mathcal{P}}[\ell_{\text{c}}(\hat{\lambda};x,y,c)]\leq\epsilon\>,$
where $\epsilon>0$ is a user-specified tolerance level representing the acceptable degree of overthinking. Naturally, smaller values of $\epsilon$ impose stricter control, prioritizing thresholds that suppress overthinking more aggressively. However, this may come at the cost of reduced performance on helpful context—a tradeoff we explore in §H. Since we observe that risks computed using the context-aware loss $\ell_{\text{c}}$ are not monotonic with respect to $\lambda$ (Fig. 9), the Learn-then-Test (LTT) framework (Angelopoulos et al., 2021) is the only viable option; hence, we use LTT as our risk-control approach for selecting our exit threshold $\hat{\lambda}$.